Data Table Molarity of Na2S2O3 0.125 M

Initial Buret Reading (Na2S2O3) 1.50 mL

Final Buret Reading (Na2S2O3) 14.35 mL 

Calculate the average molarity of commercial bleach before it was diluted


1
Expert's answer
2021-01-13T13:50:24-0500

 Molarity of Na2S2O3 0.125 M

Initial Buret Reading (Na2S2O3) 1.50 mL.


So , MilliMoles of Na2S2O3

= ( 1.50 * 0.125 ).


Final Average Concentration = 1.50 * 0.125 / 14.35 = 0.1306620 M .